# OptiMUS-0.3

### v1

@@ 5.2 Ablation Study
Table 3 shows the impact of different components, as well as choice of LLM on the performance of OptiMUS. One interesting observation is the performance drop that occurs when less capable LLMs are used instead of GPT-4-o. Optimization tasks need complicated reasoning, and we believe that smaller models are mostly optimized on simpler tasks. Nonetheless, OptiMUS-0.3 can achieve an acceptable performance using the open-source LLaMa3-70B.
We can observe that the constraint extraction error correction step is cruicial to OptiMUS-0.3’s performance. The modeling EC, however, is not as important, mainly because most of the modeling errors can later be fixed in the debugging step (while constraints can only be added in the extraction step).
The LLM feedback step marginally improves the performance of the system. The debugging step, is particularly important for harder problems, given LLMs often make a few small mistakes when they generate code.
Fig. 12 left shows the relation between the performance and the number of debugging steps for different models. In all models, the first few debugging steps increase the performance, but the gain plateaus afterwards. Fig. 12 shows the composition of different failure reasons for easy and hard problems. Clause extraction and clause modeling are the most common reason of failure.
@@ 5.3 Failure Cases
To understand its strengths and weaknesses, we analyze the most common reasons why OptiMUS fails. We categorize failure cases into the following groups (Fig. 12):
  • Missing or wrong clauses: OptiMUS generates wrong constraints or objective (e.g., price $\geq$ 0 where price is a parameter), or fails to extract all of the constraints from the description.
  • Incorrect model: OptiMUS tackles the problem with an incorrect mathematical model (e.g., defining binary variables for visiting cities instead of links in TSP).
  • Coding error: OptiMUS does not generate error-free code even after debugging. Coding errors often occur when the LLM is confused by the language used (e.g. a runtime error because of accessing parameter array indices that do not exist).
Understanding and interpreting the problems is also challenging for LLMs, resulting in formulations with missing constraints and wrong constraints. Fine-tuning might improve the performance of LLMs on this task, and is an important direction for future research.
Table 2: OptiMUS-0.3 outperforms the other methods
@@
OptiMUS-0.3 88.6% 73.7%
Table 3: Ablation studies’ results on OptiMUS-0.3
@@
ExtractionModelingCoding0020202020404040406060606080808080100100100100Failure TypeRate(%)EasyHard
Figure 12: Left) More debugging iterations improves the performance. Right) For harder problems, most failures happen during clause extraction. For easier problems, most failures happen during the coding step.
Table 4: Error correction methods can find and fix a large fraction of errors in constraint extraction (left) and constraint modeling (right), while leaving the correct items intact most of the time (a diagonal matrix is good)
@@
5.4 RAG
Figure 13: Retrieval Augmented Generation (RAG) can improve the performance of OptiMUS-0.3
Easy Hard
OptiMUS-0.3 88.6% 73.7%
RAG on Clauses 83.8% 79.6%
RAG on Problem Description 86.7% 45.5%
RAG on Problem Labels 87.1% 53.6%
Best 88.6% 79.6%
Retrieval Augmented Generation (RAG) enhances the performance of the LLMs by augmenting the prompt with specialized data relevant to the task. It has shown success in Q&A systems that require domain-specific knowledge Lewis et al. (2020), Gao et al. (2023b), Liu et al. (2020). The RAG module in OptiMUS-0.3 aims to improve the ability of the system to correctly model objectives and constraints by retrieving examples of similar constraints and their formulations and including them in the LLM prompt. The RAG-dataset consists of the 218 problems from the dev and test sets for which OptiMUS produced a correct formulation. For a given input problem, the RAG module of OptiMUS searches for examples from among all problems in the RAG-dataset, excluding the current input problem.
We employ RAG in two phases of Algorithm 1: during the extraction of constraints and objectives (ExtractTargets, line 4), and during the formulation of constraints and objectives (FormulateClauses, line 7). In ExtractTargets, we retrieve the five problem descriptions from the dataset that have the highest cosine similarity to the OpenAI (text-embedding-3-large) embeddings of the problem descriptions. These retrieved descriptions, along with their corresponding natural language objectives and formulations, are prepended to the original prompt. In FormulateClauses, we identify the ten constraints or objectives that are most similar in terms of cosine similarity to the natural language descriptions of the current objective or constraint. We then prepend only these objectives or constraints, not the problem description, to the prompt. We also explored using similar problems in FormulateClauses, by adding three problem descriptions with natural language and formulation of objectives/constraints, by either comparing the similarity of the textual descriptions or Jaccard similarity of the problem labels from Figure 1. In Fig. 13, notice that for the easy problems, all of these RAG variants decrease performance, but for hard problems, RAG on individual clauses can improve performance substantially. This might suggest that if the problem is easy and the LLM is capable of solving it, adding RAG confuses it, but if it needs to employ a technique such as big-M, having similar examples can increase performance.
To understand how RAG affects the performance, we hand-crafted some examples. Fig. 14 (left) shows a RAG setup where we pass the constraint “No more than $20\%$ of the article prices can change” alongside a randomly chosen example of a correctly formulated constraint. The model fails to correctly formulate the intended constraint. Conversely, in Fig. 14 (right), the same constraint is provided along with an example demonstrating the big-M technique. This relevant example guides the model to correctly formulate the constraint.
Hence we see that a well-chosen RAG example can improve performance, while a poorly chosen example can degrade performance. We hypothesize that irrelevant examples can distract the LLM from focusing on the constraint at hand. We suspect that overcoming this hurdle will present a substantial challenge, and so we leave to future work further development of RAG for optimization modeling. An improved RAG might replace the default distance metric used in the RAG vector search, cosine similarity, with one tailored to the optimization context (including search that is agnostic to variable name and aware of constraint type).
Figure 14: RAG helps with the relevant examples can help OptiMUS model the constraints correctly (right), while using RAG with irrelevant examples can confuse OptiMUS (left).
6 Optimization Techniques in OptiMUS
Many of the advances in optimization solvers in the last several decades rely on the fact that solvers are faster when they can exploit particular structures within the optimization problem. One of the main tasks of an optimization expert is to identify these structures so as to choose an appropriate solver or parameter settings. However, detecting useful structures in a given optimization problem can be very challenging, since it requires both a deep understanding of the mathematical structure of the problem and the capabilities of existing solvers.
In view of this challenge, OptiMUS-v0.3 includes an optimization technique detection module, which we describe in this next section. The structure detection module is composed of two agents:
  • a structure detection agent optimization structures
  • a customized optimization coding agent
Figure 15: Structure pool in OptiMUS
6.1 Design of Structure Detector and Advanced Optimization Coding
OptiMUS maintains a pool of optimization structures which are commonly exploited in optimization software. To detect structure in an optimization problem, OptiMUS iterates through these structures and format them into a structure detection prompt. Within each prompt, the LLM is provided with the description of the structure, explained by an example illustrating how the structure should be exploited. The LLM is asked to decide whether the structure can be applied to the existing formulation. Upon identifying the appropriate structure, the formulation is adjusted to high-light the problem structure.
Currently, OptiMUS considers the following three types of structures.
Problem structure
These structures apply when MILP formulation is applied to solve problems that boil down to some more specific type. Customized solvers often exist for these problems and can dramatically improve solution speed. SAT, traveling salesman (TSP), vehicle routing (VRP) knapsack, network flow, constraint programming fall into this category.
Constraint (variable) structure
These structures are considered within an MILP model and are defined on a subset of variables or constraints. Optimization solvers can exploit these structures to improve performance Gamrath et al. (2016), and customized interface are often available. Using the interface not only reduces the complexity of (and potential for errors in) using auxiliary variables or constraints, but also informs the solver of the existence of structure that can be exploited to solve the problem faster. Typical examples of structure include Special Ordered Set (SOS) Beale and Forrest (1976), indicator variables, semi-continuous variables, and general constraints Bertsimas and Tsitsiklis (1997a).
Large-scale optimization technique and advanced optimization coding
The aforementioned structures are generally straight-forward to exploit: either changing the optimization solver or modifying a few lines of modeling code suffices. However, in large scale optimization, it is common that exploiting structure becomes highly nontrivial: in these applications, optimization solvers are often embedded into some high-level optimization framework and called as a subroutine. Typical examples including decomposition structure, column generation, and cutting plane methods. To leverage these structures, optimization coding needs to invoke advanced solver interface (for example, call-back, model attribute query and analysis), which goes far beyond modeling itself. OptiMUS implements a dedicated coding agent that adapts available modeling code to some user-specified framework.
To illustrate the performance of OptiMUS, we provide showcase examples to demonstrate the efficiency of OptiMUS in optimization structure detection.
Example: Indicator Constraint Definition An indicator constraint over a binary variable $z$ and a constraint $C$ states that $z=1\Rightarrow C$. Description A retail company wants to open new stores in a set of $K$ potential locations. A store at location $k$ must have a minimum staff of $n_{k}$ employees (to ensure adequate customer service) and a minimum inventory level of $l_{k}$ units (to avoid stockout). The maximum inventory level for store $k$ is $u_{k}$, and there is a fixed cost associated with opening each store. Objective The company aims to maximize the total expected revenue from the opened stores. Relevant Constraints If a store at location $k$ is open, it must have at least $n_{k}$ employees. Moreover, let the inventory level of store $k$ be $y_{k}$. Formulations Original: $x_{k}\in\{0,1\},~{}y_{k}\leq u_{k}x_{k},$ $z_{k}\geq l_{k}-M(1-x_{k})$ OptiMUS: $x_{k}$$\Rightarrow$ $y_{k}\leq u_{k}$ $x_{k}$$\Rightarrow$ $z_{k}\geq l_{k}$ Performance plot [Uncaptioned image]
Example: Special Ordered Set Definition Given a set of variables $\mathcal{X}=\{x_{1},\ldots,x_{n}\}$, $\{x_{1},\ldots,x_{n}\}\in\operatorname{SOS}_{1}$ if at most one element of $\mathcal{X}$ can be nonzero $\{x_{1},\ldots,x_{n}\}\in\operatorname{SOS}_{2}$ if at most two elements of $\mathcal{X}$ can be nonzero Description An airline needs to assign $N$ crew members to a set of $K$ flights for a day. Each crew member can only be assigned to at most one flight at a time, and there must be a minimum rest period of 2 hours between consecutive flights to comply with aviation regulations and ensure crew well-being. Each crew member has a maximum number of flying hours per day, set at 8 hours, to comply with aviation regulations and prevent fatigue. There are minimum and maximum layover times between flights: minimum layover of 2 hours (to allow for crew rest and flight preparations) and maximum layover of 6 hours (to optimize crew utilization and reduce idle time). Assigning a crew member to a specific flight has some cost. Objective Minimize the total crew cost while ensuring that each flight has the required number of qualified crew members, satisfying the rest period constraints, maximum flying hours per crew member, and layover time constraints. Relevant Constraints Each crew member can work on only one flight. Let $x_{ij}\in\{0,1\}$ denote whether crew $i$ works on flight $j$ and $T_{ij}$ denote the working time of crew $i$ on flight $j$. Formulations Original: $\textstyle\sum_{j=1}^{K}x_{ij}\leq 1,\quad$ $\forall i\in[N],$ $T_{ij}\leq 24x_{ij},\quad$ $\forall i\in[N],j\in[K]$ OptiMUS: $\{T_{i1},\ldots,T_{iK}\}\in\operatorname{SOS}_{1}$ Performance plot [Uncaptioned image]
6.1.1 Large-scale column/constraint generation
Consider a general optimization problem
$\begin{array}[]{ll}\mbox{minimize}&c(x)\\ \text{subject to}&x\in\mathcal{X}_{k},~{}k=1,\ldots,m\end{array}$ ($\mathcal{P}$)
with variable $x\in\mathbb{R}^{n}$ in which either $m$ or $n$ is large. Column and constraint generation methods, respectively, allow for scalable solutions to such problems. For example, constraint generation for $m\gg n$ proceeds by first restricting optimization problem to a subset $\mathcal{W}\subseteq[m]$ of constraints, often chosen at random: define $x_{\mathcal{W}}^{\star}$ to be the solution of the problem
$\begin{array}[]{ll}\mbox{minimize}&c(x)\\ \mbox{subject to}&x\in\mathcal{X}_{k},~{}k\in\mathcal{W}\end{array}$ ($\mathcal{P}_{\mathcal{W}}$)
with variable $x\in\mathbb{R}^{n}$. When the number of constraints $|\mathcal{W}|$ is small, the resulting problem ($\mathcal{P}_{\mathcal{W}}$) can be solved more efficiently. Given the solution $x_{\mathcal{W}}^{\star}$ to ($\mathcal{P}_{\mathcal{W}}$), constraint generation solves a pricing problem to find the most violated constraint
[…]
This constraint $j_{\mathcal{W}}$ is added to the set $\mathcal{W}$ and the algorithm continues until convergence. As the algorithm continues until all constraints are satisfied, the solution to the final restricted master problem ($\mathcal{P}_{\mathcal{W}}$) is also optimal for the original problem ($\mathcal{P}$).
We provide an example of constraint generation in OptiMUS in Fig. 16 on a security-constrained unit commitment problem. For details of the problem see the appendix.
Figure 16: Left: constraint generation prompt. Right: performance plot of solving to 5% gap.
@@ 7 Conclusion
How can we leverage LLMs to achieve complex goals? This paper interrogates this question in the domain of optimization and showcases the importance of modular structure. We develop OptiMUS-0.3, a modular LLM-based agent designed to formulate and solve optimization problems from natural language descriptions. Our research serves as a proof-of-concept, illustrating the potential for automating various stages of the optimization process by combining LLMs with traditional solvers. To showcase the performance of OptiMUS-0.3, we released NLP4LP, a comprehensive dataset of long and challenging optimization problems to demonstrate the efficacy of the techniques implemented within OptiMUS-0.3. OptiMUS-0.3 achieves SOTA performance across all existing datasets, and scales to problems with large amounts of data and long descriptions.
Real-world optimization problems are often complex and multifaceted. Developing LLM-based solutions for these problems requires domain-specific considerations, including integrating existing optimization techniques to leverage problem structure. We are at the early stages of this research, but anticipate significant developments that will enable these systems to address more complex, industrial-level problems. It is interesting to notice that the challenge of using AI for an applied domain is much larger in safety-critical domains such as self-driving, which demand extremely high accuracy, than in domains where AI can function as an assistant and where answers are easy to check, as in theorem-proving or optimization. Here, AI systems with moderate accuracy can still usefully augment human effort.
@@
The development of OptiMUS opens up several promising avenues for future research in AI-assisted optimization modeling. Enhancing LLMs’ capacity to interpret nuanced or ambiguous problem descriptions could improve its robustness and applicability, potentially through specialized language models fine-tuned on optimization terminology. An interesting area of exploration could be the combination of structured pipelines and more flexible agent-based systems, shedding light on their relative strengths and integration potential. The development of mechanisms for efficiently integrating user feedback and iteratively refining models presents another intriguing research direction. Exploring the applicability of OptiMUS’s modular LLM structure to other complex problem-solving tasks beyond optimization might yield insights into building more capable AI systems for a range of applications. Finally, the creation of more comprehensive benchmarks for assessing AI-assisted optimization modeling tools could drive progress in this field.
Appendix: SCUC problem
Description.
The Security Constrained Unit Commitment (SCUC) is a critical optimization problem in the operation of electrical power systems. Its goal is to schedule generation units in a cost-effective manner while ensuring reliable operation over a specified time period (typically 24 hours), considering the system’s physical and operational constraints. This appendix presents a detailed description of the SCUC problem, including its objectives and constraints:
Objective.
The objective of the SCUC problem is to minimize the total operational cost of the power system. This cost generally includes:
  • Fuel Cost: The cost associated with consuming fuel to generate electricity.
  • Start-up and Shut-down Cost: Cost incurred from starting up or shutting down generation units.
  • Emission Cost: Cost related to the emissions produced by the generating units, if applicable.
Constraints.
The SCUC must satisfy several constraints to ensure the safe and reliable operation of the power system:
  • Power Balance Constraint: Ensures that the total power generation meets the total demand plus system losses at every interval. Mathematically, this is expressed as the sum of the outputs of all online generators being equal to the sum of the demand and the transmission losses.
  • Minimum and Maximum Output Limits: Each generator has a minimum and maximum generation capacity when it is online. The SCUC ensures that the output of each unit stays within these bounds.
  • Ramp-up and Ramp-down Limits: These limits specify the maximum rate at which a generator can increase or decrease its output. They are critical for handling load changes throughout the day.
  • Unit Start-up and Shut-down Constraints: These constraints handle the logistics of turning units on or off. Start-up constraints may include minimum down times (the minimum time a unit must remain off before it can be restarted) and minimum up times (the minimum time a unit must remain on once started).
  • Minimum Up and Down Time Constraints: Ensures that once a generator is turned on, it stays on for at least its minimum up time, and similarly, once turned off, it remains off for at least its minimum down time. These constraints are crucial for the mechanical integrity of the generation units.
  • Reserve Requirements: The system operator must ensure that sufficient spinning and non-spinning reserves are available. These reserves are needed to handle sudden increases in demand or unexpected generator failures.
Appendix: Checking formulation equivalence
To evaluate the correctness of the formulation of a given optimization problem, OptiMUS evaluates the formulation on a numerical example and checks that the optimal value matches the ground truth and that all constraints are satisfied. This approach runs the risk that an incorrect formulation might return the correct solution on the particular problem data chosen, but not for different problem data. An alternative approach would be to compare the formulation against a reference ground-truth mathematical formulation, and evaluate whether the formulations are the same. This alternative approach has the benefit that it is independent of numerical data, and indeed would not require the availability of any problem data. This benefit is important for complex and realistic industrial problems, for which true problem data may be proprietary and feasible problem data may be difficult to generate.
Unfortunately, there are several serious obstacles to comparing the formulation of a problem against a reference formulation. First, for most reasonable notions of equivalence, two formulations can both be correct even if they are not equivalent. Worse, the problem is NP-hard: in its general form, the problem of checking the equivalence of formulations can be reduced to (a general form of) the well-known Graph Isomorphism Problem (GIP). In the discussion that follows, we restrict our attention to feasibility problems for simplicity; the extension to more general optimization problems is similar.
Formulation equivalence
Two mathematical formulations are considered equivalent if there exists a one-to-one correspondence between their variables and constraints in which the coefficients of corresponding pairs of variables are equal in every corresponding pair of constraints. In other words, the coefficients of each variable, whether symbolic or numeric, must align under the mapping. Note that it is possible for two formulations to describe the same feasible set without such a correspondence, in which case they are not considered equivalent under this strict definition.
Reducing formulation equivalence to generalized graph isomorphism
To compare two formulations, we first represent each formulation as a bipartite graph $G$ where variables and constraints form separate sets of nodes. A variable node is linked to a constraint node by an edge if the variable appears in the constraint, with the edge labeled by the variable’s coefficient in the constraint. This unique graph representation allows us to show the equivalence of two formulations by establishing an isomorphism between their respective graphs.
Given two formulations, $\mathcal{F}$ and $\mathcal{F}^{\prime}$, with differing variables and constraints, we construct their corresponding graphs $\mathcal{G}$ and $\mathcal{G}^{\prime}$. The challenge then is to find a correspondence between the graphs’ nodes that respects the connectivity and coefficient labels, effectively determining graph isomorphism between $\mathcal{G}$ and $\mathcal{G}^{\prime}$.
Reducing graph isomorphism to formulation equivalence
Consider two graphs $\mathcal{G}_{1}=(V_{1},E_{1})$ and $\mathcal{G}_{2}=(V_{2},E_{2})$. To check for isomorphism, we construct corresponding mathematical formulations, $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$.
Constructing mathematical formulations
  1. 1.
    Variables and constraints: For each vertex $v$ in $V_{i}$, define a variable $x_{v}$. For each edge $(u,v)\in E_{i}$, create a constraint $x_{u}+x_{v}=1$.
  2. 2.
    Formulation: Each formulation $\mathcal{F}_{i}$ includes:
    • Variables: $\{x_{v}\mid v\in V_{i}\}$
    • Constraints: $\{x_{u}+x_{v}=1\mid(u,v)\in E_{i}\}$
An algorithm that finds a bijective mapping between the variables and constraints of $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ effectively establishes a one-to-one correspondence between the vertices and edges of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$.
This mapping ensures:
  • Each variable in $\mathcal{F}_{1}$ maps to a unique variable in $\mathcal{F}_{2}$, corresponding to a vertex in $\mathcal{G}_{2}$.
  • Each constraint in $\mathcal{F}_{1}$ maps to a unique constraint in $\mathcal{F}_{2}$, corresponding to an edge in $\mathcal{G}_{2}$.
Thus, verifying the equivalence of $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ is equivalent to determining if $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are isomorphic, showing that this problem is as computationally complex as the Graph Isomorphism Problem.
Concluding the reduction
This approach effectively reduces the problem of checking graph isomorphism between $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ to verifying the equivalence of the mathematical formulations $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$. This reduction not only underscores the computational complexity of formulation equivalence but also classifies it within the same complexity category as the (NP-complete) Graph Isomorphism Problem.

### v2

@@ 5.2 Ablation Study
Table 3 shows the impact of the choice of LLM on the performance of OptiMUS. We evaluate the OptiMUS system with both a leading closed-source LLM, GPT-4o, as well as a weaker open-source alternative, LLaMa-3.1-70B-Instruct (Meta AI 2024). As expected, the performance of the OptiMUS-0.3 framework degrades when less capable LLMs, such as LLaMa-3.1-70B-Instruct, are used instead of GPT-4o, especially on hard modeling tasks. Optimization tasks need complex reasoning, and smaller models may be better suited for simpler tasks. However, the OptiMUS-0.3 framework with LLaMa-3.1-70B-Instruct outperforms GPT-4o on with previous state-of-the-art prompting strategies, including Reflexion, on easy problems. This result highlights that the modular design of OptiMUS-0.3 not only allows it to outperform other prompting strategies using the same model, but can also makes the system more robust to the quality of the underlying LLM, allowing weaker cheaper models to match the performance of stronger models on simple optimization modeling tasks.
Table 3 also shows the impact of different components of the OptiMUS framework on its performance. Debugging (i.e., running error correction on the output of the Coding stage) has the largest impact on the performance of OptiMUS, increasing its accuracy on hard problems by almost 50%. The debugging step, is particularly important for harder problems, given LLMs often make a few small mistakes when they generate code. Error correction also leads to modest improvements in the performance of OptiMUS. Running error correction after extracting clauses is particularly important for difficult problems. The LLM feedback step marginally improves the performance of the system. Note that we do not include the impact of the structure detection agent and advanced optimization coding agent as they do not impact the accuracy of the framework, but just the efficiency (i.e., runtime). For examples showcasing the impact of these components, see Appendix 11.
Fig. 7 (left) shows the relation between the performance and the number of debugging steps for different GPT-4o and LLaMa-3.1-70B-Instruct. In both models, the first few debugging steps improve performance, but further steps do not result in improvement.
@@ 5.3 Failure Cases
We analyze the most common reasons why OptiMUS fails, categorizing failure cases into the following groups (Fig. 7):
  • Extraction errors: OptiMUS extracts the wrong natural language constraints or objective (e.g., price is non-negative where price is a parameter), or fails to extract all of the constraints from the description.
  • Formulation errors: OptiMUS formulates a mathematical model that does not match the natural language constraint (e.g., it uses the wrong parameter to lower-bound a variable).
  • Coding errors: OptiMUS does not always generate error-free code even after debugging. Coding errors most often occur when the LLM is confused by the language used (e.g. a runtime error due to accessing indices that do not exist in a parameter array).
As shown in Fig. 7, OptiMUS successfully extracts almost all the correct clauses for the easy dataset. This task is easy on the easy dataset because the easy instances are all LPs and involve only scalar values. However, for the hard instances — comprising MILP problems and multi-dimensional variables — parameter extraction and modeling are considerably more challenging than coding the resulting model. Note that the coding error rate is lower for the hard dataset because OptiMUS sometimes fails in modeling or formulating hard clauses, producing clauses that cannot be coded, which we do not include in the coding statistics.
Table 2: OptiMUS-0.3 outperforms the other methods with GPT-4o
@@
OptiMUS-0.3 86.6% 73.7%
Table 3: Ablation studies on OptiMUS-0.3
@@
ExtractionModelingCoding00101010102020202030303030404040405050505060606060707070708080808090909090100100100100Failure TypeError rate (%)EasyHard
Figure 7: Left) Further debugging iterations improve performance. Right) For harder problems, most failures arise from clause extraction mistakes. For easier problems, most failures are due to coding errors.
Table 4: Error correction methods can find and fix a large fraction of errors in constraint extraction (left) and constraint modeling (right), without modifying most correct items. (Perfect performance is diagonal.)
@@
6 Web Application
As discussed in 1, the OptiMUS project aims to 1) help optimization experts develop and maintain their models with ease and 2) empower domain experts to make better decisions faster and reduce costs. To further these goals, we designed and built a user-centric web app with an intuitive interface that enables users to interact seamlessly with the system and leverage the advantages of large language models (LLMs). The webapp allows users to follow the overall process of modeling an optimization model including extracting parameters and clauses, modeling each clause, and finally coding it up. Importantly, users can correct any output of the framework, including natural language clauses, model, and code, to provide better supervision over the entire process. By allowing users to observe and provide input throughout each step, the web app brings significant speed and convenience to the modeling process while reducing the risk of errors. Figure 8 shows a sample model in the Coding stage. Further details of all the different stages and features incorporated in the webapp can be found in Appendix 14. The webapp is publicly available at https://optimus-solver.com/.
Figure 8: Sample view of the OptiMUS webapp during the coding phase in which individual clauses are translated to python code. Users have the ability to see the confidence scores, correct the code directly in the panes to the right, or ask OptiMUS to regenerate code.
@@ 7 Conclusion
How can we leverage LLMs to achieve complex goals? This paper interrogates this question in the domain of optimization and showcases the importance of modular structure. We develop OptiMUS-0.3, a modular LLM-based agent designed to formulate and solve optimization problems from natural language descriptions. Our research illustrates the potential to automate aspects of the optimization process by combining LLMs with traditional solvers. To assess the performance of OptiMUS-0.3, we released NLP4LP, a comprehensive dataset of long and challenging optimization problems. OptiMUS-0.3 achieves SOTA performance and scales to problems with large amounts of data and long descriptions.
Real-world optimization problems are often complex and multifaceted. Developing LLM-based solutions for these problems requires domain-specific considerations, including integrating existing optimization techniques to leverage problem structure. We are at the early stages of this research, but anticipate significant developments that will enable these systems to address more complex, industrial-level problems. It is interesting to notice that the challenge of using AI for an applied domain is much larger in safety-critical domains such as self-driving, which demand extremely high accuracy, than in domains where AI can function as an assistant and where answers can be checked by humans, as in theorem-proving or optimization. Here, AI systems with moderate accuracy can still usefully augment human effort.
@@
The OptiMUS system presented here serves as a prototype for more powerful AI-assisted optimization modeling. Much more work remains to make such a system reliable and to facilitate the use of classical optimization tools for domain users who know little about optimization. While some improvements can be expected as language models grow more powerful, we expect that substantial breakthroughs by the optimization community will be required to address many of the following challenges:
  • Reliability: Optimization algorithms generally offer guaranteed reliability. How can a system like OptiMUS reliably solve large-scale problems using unreliable LLMs that suffer hallucinations, overconfidence, and instability? What kinds of guarantees can be made, and how expensive might these be? How should developers of such a system assess whether the system has succeeded in producing a correct model or not?
  • Trust: A system like OptiMUS should help domain experts understand the model that was selected and decide whether it meets their needs. How should the system communicate a model back to the human user to facilitate effective feedback about the deficiencies in the model, or trust in a model that is correct? How should the system integrate this feedback? How can the system developer assess the effectiveness of this feedback and improve the system to enable quicker and easier feedback cycles?
  • Ambiguity: Most optimization problems specified in natural language do not correspond unambiguously to a single optimization model. Rather, the process of formalizing the objective and constraints reduces the vagueness so that it is possible to distinguish a solution from any other proposal. A system like OptiMUS should help decision-makers think through their goals (which may be vague) to choose a model formulation by clarifying implicit requirements, eliciting the relative importance of different objectives, and searching internal or external databases for relevant data.
  • Fast Solvers: Rapid decisions require a fast solver: possibly a well-chosen solver for the problem, or possibly even a custom solver written for the problem. A system like OptiMUS should be able to choose between available heuristics and parameters for provably optimal solvers using all information available about the problem to determine a solution method that meets runtime and accuracy requirements, or even to write a custom solver when it is warranted. To choose or write the best solver, we expect that the system will likely make use of natural-language information about the problem setting (using an LLM) as well as structured information about problem parameters (using a graph neural network).
  • Larger datasets. All of these research directions depend on the availability of appropriate datasets to develop, hone, and test new ideas. Most of these directions require a dataset of optimization problems consisting of a natural-language description coupled with problem data and solution code. While there are many excellent datasets for developing optimization algorithms, these are generally not associated with natural-language descriptions; and problems associated with natural language descriptions are generally either small (e.g., textbook problems) or lack associated problem data and solutions (e.g., published papers on applications of MIPs) as these elements are often proprietary. Semi-synthetic approaches to generating larger datasets are promising (Tang et al. 2024), but creating MIP problems that are large-scale, realistic, and feasible is a challenge. Curating a larger dataset of optimization problems could be a key enabler driving further research and development in this area.
  • Beyond MILPs. OptiMUS currently uses gurobipy as a modeling language. We made this choice as the wealth of online examples allows LLMs to successfully model many problems in gurobipy. However, many optimization problems are best expressed in a different framework: for example, Concorde for traveling salesman problems (Cook et al. 2011), cvxpy for control problems (Diamond and Boyd 2016), or minizinc for constraint programming problems (Nethercote et al. 2007). Integrating these alternative modeling languages and identifying, for a given problem, which modeling language to use (or whether the problem is likely to be successfully modeled in the OptiMUS framework at all) is an important future direction.